you bought a house for R90 000. the value of your house appreciate at 10% per year. how much is your house worth after 1 year and 3 years?


1
Expert's answer
2020-10-04T17:50:43-0400

value of the house after 1 year="initial value\times(1+\frac{r}{100})^n"

"= 90000\times(1+\frac{10}{100})^1"

"= 90000\times1.1"

"=R99000"


value of the house after 3 years ="initial value \times(1+\frac{r}{100})^n"

"= 90000\times(1+\frac{10}{100})^3"

"= 90000\times1.331"

"=R11979"